# Generation of a scalar vortex in a rotational frame

### v1

@@ Abstract
We consider generation from the vacuum of a scalar charged field in a rigidly rotating frame. Adding an external magnetic field opens the way to Bose condensation of the field. This phenomenon has been studied for external uniform magnetic field occupying the whole volume of the uniformly rotating cylindrical system of finite radius $R$ with a Dirichlet boundary condition imposed on it. Besides continuing this study, we consider the field formed by a flux tube of small radius. We find numerical solutions of the Ginzburg-Pitaevskii equation for the charged scalar field, the critical rotation frequencies, the mean radii and the condensate energies, and compare them with those found in a linearization scheme and with approximate analytical solutions. We show that for the same input parameters the energy of the condensate in the case of the flux tube is lower than in the case of uniform magnetic field in the whole cylinder.
@@ 2.1 Resting and rotating frames
${T^{\mu}}_{\nu}=\frac{\partial x_{R}^{\mu}}{\partial x^{\nu}}={\left(\begin{array}[]{cc}1&-\vec{v}_{R}^{\top}\\ 0&\bm{D}^{\top}(\Omega t)\end{array}\right)^{\mu}}_{\nu},$ (25)
@@ 2.2 Lagrangian and field equation in the rotational frame
where $-e$ is the charge of the electron, $e^{2}\approx 1/137$, $\partial_{\mu}$ in the rotational frame is determined in (36). Although the most physically meaningful case is that of the pseudoscalar charged pion field, for definiteness let us further speak about the scalar field. In the rotational frame it is a function of coordinates, $\phi(t,\vec{x})$, which we suppress temporarily for simplicity of the formulas. The derivatives are partial ones for the time-space part (in the Lorentz indices); since $\phi$ is a scalar field no Christoffel symbols enter. Below, since the metric (26) is flat, we need neither $\sqrt{-g}$ nor the curvature $R$. We consider the field inside a cylinder of some radius $R$, fulfilling the causality condition $\Omega R\leq 1$ and must thereby impose the boundary condition. We take the Dirichlet condition, $\phi(\vec{x})=0$ at $r=R$. With this condition, surface terms will not appear when integrating by parts in the Lagrangian. Now, the Lagrangian density (43) can be rewritten as
@@
In the following, dealing with a static magnetic background field $\vec{A}_{R}$ we put $A_{R,0}=0$ in (10) and (39). Then we are able to focus on the case of the stationary condensate field $\phi(t,\vec{x})\propto e^{-i\omega t}\phi(\vec{x})$ occurring at the corresponding conditions in the rotational frame. The energy density corresponding to the Lagrangian density is ${\cal{E}}=\omega n-{\cal L}$, where $n=\partial{\cal L}/\partial\omega$ is the particle density. With increasing rotation speed $\Omega$, the ground state energy of the boson $\omega$ first reaches zero, $\omega=0$, and then dives further down to $-m$. Then in the passive rotation problem bosons of one sign of charge can be produced occupying the level $\omega=-m$ in a tunneling process from the lower to the upper continuum and anti-bosons go off to infinity. To be careful, note that in a short-range electric potential well at $mR\gg 1$ the anti-particle energy level appears near the boundary of the lower continuum, at $\omega\simeq-m+O(1/R)$, see [1]. In the presence of the empty rotating “physical” cylinder, instability may occur already when the ground state single-particle energy level reaches zero; then the reactions with production of scalar particles and antiparticles on the surface of the cylinder become possible. Antiparticles are absorbed by the surface of the cylinder. Particles form the classical field $\phi(\vec{x})$ inside the cylinder. This mechanism of production of the field $\phi$ was discussed in [18, 19]. For $\omega>0$ the number of particles, $n$, calculated within the single-particle problem is non-negative, see [20]. The minimum of $\omega n$ corresponds to $\omega=0$. In this case when the single-particle ground state energy level, $\omega$, reaches zero a new vacuum state can be occupied, characterized by the classical static ($\omega=0$) field $\phi$, which is described by the GP-like equation. The value $\omega=0$ plays the role of the chemical potential for the order parameter $\phi$, see [2], [3]. Equation (46) turns then into
@@ 3.1 Notations for the background field
3.2 Ginzburg-Pitaevskii like equation for the condensate
@@ 3.4.1 Solution of a linearized equation
We should recall that in our case as input parameters we use the values $R,B_{0},\Omega$ and for infinitely thin flux tube the values $R$ and $B_{0}$ enter only through the value of the magnetic flux, $\delta_{\Phi}$. As we shall see below, the value of the angular momentum $l$ in (65) corresponding to the minimal energy $E_{GP}$ is expressed via mentioned input parameters and differs only a little from $\delta_{\Phi}$ and whether in the energy minimum $l-\delta_{\Phi}>0$ or $<0$ is not determined since $\nu$ enters GP-like equation (64) only as $\nu^{2}$.
@@
The relation (70) is fulfilled even for quite large ${\tilde{\epsilon}}\lesssim 4.8$. The energy per unit length, $E_{0}^{lin}$, following from (63) with $\lambda=1$, is shown in figure 1 as function of $\Omega R$. Also we show the dependence by $\Omega R=F({\tilde{\epsilon}})$ following (68). As we see from the figure 1, the energy density $E^{lin}/(\pi R^{2})$ coincides well with the exact one for ${\tilde{\epsilon}}\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}1$ that corresponds to $\Omega R\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}0.29$, and as expected, in the whole region we have $E^{lin}>E_{0,GP}$. The first excited state solution arises from the root $j_{\nu,2}$ in the solution (67). It appears with increasing $\Omega$ only at $\Omega R\simeq 0.56$, ${\tilde{\epsilon}}\simeq 2.61$, and therefore it is not shown in figure 1. In the inset we show parameters $\Omega$ and ${\tilde{\epsilon}}$ as functions of $\Omega R$. They are very close to each other except the vicinity of the critical point.
@@ 3.4.3 Approximate analytical solution of the GP-like equation
$\phi_{appr}(r)=\frac{\tilde{\Omega}}{\sqrt{\lambda}}\tanh\left(\frac{R-r}{\sqrt{2}}\tilde{\Omega}\right).$ (75)
@@
[…]
The value $\langle r\rangle$ increases with $\tilde{\Omega}$ following the peak of the function $r\phi(r)^{2}$. As we see, in the case of the infinitely thin flux tube the mean radius changes in a rather narrow interval, $0.70.5<\langle r\rangle/R<$ 0.7, since the solution $\phi(r)$ is smooth and becomes closer to a step-function with increasing $\Omega R$.
@@
[…]
@@
[…]
@@ 3.5.1 Linearized equation
In general $\alpha$ is not necessarily an integer number. Nevertheless, a strikingly simple way to get information from the Kummer function is to terminate the series expansion (85) by taking for $-\alpha$ a natural number, which was used in [16]. In the latter case the minimal energy corresponds to $\alpha=-1$. If one took $\alpha=-2$ or smaller, one could find the relations (more than one and non-linear) between the flux and the orbital number, for which the boundary conditions are fulfilled. All these have $\delta_{\Phi}>l$. So, taking $\alpha=-1$ and $\rho=\beta$ in (85), we arrive at a solution of the equation for the boundary condition, which is valid with $\beta$ from (84) and $\delta_{\Phi}$ from (54) for
@@ 3.5.2 Numerical solutions of GP-like equation
In figure 6, the solutions of the linearized and the exact equations are shown for $l\neq\delta_{\Phi}$ at $\Omega R<1$. In both cases the solutions are rather close to one another, more so for smaller $\Omega R$, $\tilde{\Omega}$, ${\tilde{\epsilon}}$, as expected. The solutions have a Gaussian-like form and the maximum of $\phi$ and mean radius are shifted to $r$ closer to $R$ and the plateau which we observed in the case of a thin flux tube is not formed. This is because the asymptotic solution at $r\to 0$ behaves as $\phi\propto r^{l}$ and here $l\gg 1$. In the case of the thin flux tube we had $\phi\propto r^{\nu}$ and considered not large values of $\nu$ since the minimal energy corresponds then to $\nu=0$. In figure 7 we demonstrate that a plateau would be nevertheless formed also in the case of the field in the full cylinder but in the not causal region, for $\Omega R>1$.
@@ 4 Conclusion
We started with the derivation of the Ginzburg-Pitaevskii-like equation for the charged scalar (pseudoscalar) mean field in the uniformly rotating frame in the presence of the external electromagnetic background field introduced initially in the resting frame. We explicitly performed the transition from variables in the resting frame to the rotating frame used in the sense of a passive transform. We mention that for the charged scalar field the Klein-Gordon equation in the rotating frame is the same as in the local flat frame, which is also used for this kind of problems. The causality condition requires that $\Omega R<1$, which is ensured by imposing the Dirichlet boundary condition at $r=R$. It raises the question of to what extent this circumstance influences the dynamics of the scalar field. As can be seen in figure 6, where we showed the behavior of the scalar field in the rotating frame in the presence of the external uniform magnetic field, the condensate field is concentrated near the surface of the “mathematical” cylinder, which implies a large influence. Thereby we may conclude that passive rotation of the frame and active rotation of the empty “physical” cylinder with a non-transparent wall represent, in a sense, different cases. Although in the former case interpretation causes questions, the latter case allows for a clearer physical interpretation. Thereby we focused on the latter case. Thus we related the metric to the rotating empty cylinder and considered the case when the flux tube (solenoid) of radius $R_{s}\leq R$ generating the uniform magnetic field for $r<R_{s}$, see eq. (52), is at rest in the frame associated with the rotating cylinder of radius $R$ (rotating frame). The aim was to describe the response of the scalar field vacuum to such a magnetic field in the rotating frame.
@@
In the foregoing section, to be specific, we considered two configurations of the magnetic field: an infinitely thin flux tube (the solenoid of radius $R_{s}\ll R$ generating uniform magnetic field inside it) and a field homogeneous in the whole cylinder (the solenoid of radius $R$). The case of the uniform magnetic field in the whole volume of the rotating cylinder was previously studied in [16, 17, 20]. The case of a thin flux tube in the rotating frame was not studied before, to the best of our knowledge. In both cases we found a numerical solution of the linearized equation, as an approximation, and a numerical solution of the full non-linear Ginzburg-Pitaevskii-like (GP-like) equation, as well as found approximate analytical solutions, where possible. For the solution of the linearized equation, in the case of the infinitely thin flux tube, the condensate energy is expressed in terms of zeros of the Bessel function. For the case of the uniform magnetic field in the cylinder one needs to find a zero of the Kummer function, which problem is more delicate, especially for large values of parameters like the rotation frequency. However, for some special relations between magnetic flux and orbital momentum, e.g., when the magnetic flux $\delta_{\Phi}$ and the orbital momentum $l$ are connected as $\delta_{\Phi}=l+1$, an explicit solution exists. In the numerical approach, we adopted a relaxation method and a 4th order Runge-Kutta approach. Moreover, for the thin flux tube in the case $\delta_{\Phi}=l$ we found the approximate analytical solution of the GP-like equation. The latter case is of special interest, since we showed that the energy of the produced condensate takes a minimum precisely for $l=\delta_{\phi}$, see figure 3.

### v2

@@ Abstract
We consider generation from the vacuum of a scalar charged field in a rigidly rotating frame. Adding an external magnetic field opens the way to Bose condensation of the field. This phenomenon has been studied for external uniform magnetic field occupying the whole volume of the uniformly rotating cylindrical system of finite radius $R$ with a Dirichlet boundary condition imposed on it. Besides continuing this study, we consider the field formed by a flux tube of small radius. We find numerical solutions of the Gross-Pitaevskii equation for the charged scalar field, the critical rotation frequencies, the mean radii and the condensate energies, and compare them with those found in a linearization scheme and with approximate analytical solutions. We show that for the same input parameters the energy of the condensate in the case of the flux tube is lower than in the case of uniform magnetic field in the whole cylinder.
@@ 2.1 Resting and rotating frames
${T^{\mu}}_{\nu}=\frac{\partial x_{R}^{\mu}}{\partial x^{\nu}}={\left(\begin{array}[]{cc}1&0\\ -\vec{v}_{R}&\bm{D}(\Omega t)\end{array}\right)^{\mu}}_{\nu},$ (25)
@@ 2.2 Lagrangian and field equation in the rotational frame
where e-e is the charge of the electron, $e^{2}\approx 1/137$, $\partial_{\mu}$ in the rotational frame is determined in (36). Although the most physically meaningful case is that of the pseudoscalar charged pion field, for definiteness let us further speak about the scalar field. In the rotational frame it is a function of coordinates, $\phi(t,\vec{x})$, which we suppress temporarily for simplicity of the formulas. The derivatives are partial ones for the time-space part (in the Lorentz indices); since $\phi$ is a scalar field no Christoffel symbols enter. Below, since the metric (26) is flat, we need neither $\sqrt{-g}$ nor the curvature $R$. We consider the field inside a cylinder of some radius $R$, fulfilling the causality condition $\Omega R\leq 1$ and must thereby impose the boundary condition. We take the Dirichlet condition, $\phi(\vec{x})=0$ at $r=R$. With this condition, surface terms will not appear when integrating by parts in the Lagrangian. Now, the Lagrangian density (43) can be rewritten as
@@
[…]
@@ 3.1 Notations for the background field
3.2 Gross-Pitaevskii like equation for the condensate
@@ 3.4.1 Solution of a linearized equation
We should recall that in our case as input parameters we use the values $R,B_{0},\Omega$ and for infinitely thin flux tube the values $R$ and $B_{0}$ enter only through the value of the magnetic flux, $\delta_{\Phi}$. As we shall see below, the value of the angular momentum $l$ in (65) corresponding to the minimal energy $E_{GP}$ is expressed via mentioned input parameters and differs only a little from $\delta_{\Phi}$ and whether in the energy minimum $l-\delta_{\Phi}>0$ or <0<0 is not determined since $\nu$ enters GP-like equation (64) only as $\nu^{2}$.
@@
[…]
@@ 3.4.3 Approximate analytical solution of the GP-like equation
[…]
@@
$\langle r\rangle=\frac{\int dx\,r\,\phi(\vec{x})^{2}}{\int dx\,\phi(\vec{x})^{2}}.$ (76)
The value $\langle r\rangle$ increases with $\tilde{\Omega}$ following the peak of the function $r\phi(r)^{2}$. As we see, in the case of the infinitely thin flux tube the mean radius changes in a rather narrow interval, $0.5<\langle r\rangle/R<0.7$, since the solution $\phi(r)$ is smooth and becomes closer to a step-function with increasing $\Omega R$.
@@
$E_{GP}=-\frac{\pi R^{2}\tilde{\Omega}^{4}}{\lambda}\int_{0}^{\tilde{\Omega}R}dy\,y\ \chi^{4}(y)\,.$ (77)
@@
To arrive at this result we used that $\int_{0}^{y_{0}}y\mbox{th}^{4}ydy\simeq y_{0}^{2}/2-4y_{0}/3$ for $y_{0}\gg 1$.
@@ 3.5.1 Linearized equation
In general $\alpha$ is not necessarily an integer number. Nevertheless, a strikingly simple way to get information from the Kummer function is to terminate the series expansion (85) by taking for $\alpha$ a natural number, which was used in [16]. In the latter case the minimal energy corresponds to $\alpha=-1$. If one took $\alpha=-2$ or smaller, one could find the relations (more than one and non-linear) between the flux and the orbital number, for which the boundary conditions are fulfilled. All these have $\delta_{\Phi}>l$. So, taking $\alpha=-1$ and $\rho=\beta$ in (85), we arrive at a solution of the equation for the boundary condition, which is valid with $\beta$ from (84) and $\delta_{\Phi}$ from (54) for
@@ 3.5.2 Numerical solutions of GP-like equation
In figure 6, the solutions of the linearized and the exact equations are shown for $l\neq\delta_{\Phi}$ at $\Omega R<1$. In both cases the solutions are rather close to one another, more so for smaller $\Omega R$, $\tilde{\Omega}$, $\tilde{\epsilon}$}, as expected. The solutions have a Gaussian-like form and the maximum of $\phi$ and mean radius are shifted to $r$ closer to $R$ and the plateau which we observed in the case of a thin flux tube is not formed. This is because the asymptotic solution at $r\to 0$ behaves as $\phi\propto r^{l}$ and here $l\gg 1$. In the case of the thin flux tube we had $\phi\propto r^{\nu}$ and considered not large values of $\nu$ since the minimal energy corresponds then to $\nu=0$. In figure 7 we demonstrate that a plateau would be nevertheless formed also in the case of the field in the full cylinder, however not in the causal region, but beyond, in the region $\Omega R>1$, which is considered unphysical in the context of this paper.
@@ 4 Conclusion
We started with the derivation of the Gross-Pitaevskii-like equation for the charged scalar (pseudoscalar) mean field in the uniformly rotating frame in the presence of the external electromagnetic background field introduced initially in the resting frame. We explicitly performed the transition from variables in the resting frame to the rotating frame used in the sense of a passive transform. We mention that for the charged scalar field the Klein-Gordon equation in the rotating frame is the same as in the local flat frame, which is also used for this kind of problems. The causality condition requires that $\Omega R<1$, which is ensured by imposing the Dirichlet boundary condition at $r=R$. It raises the question of to what extent this circumstance influences the dynamics of the scalar field. As can be seen in figure 6, where we showed the behavior of the scalar field in the rotating frame in the presence of the external uniform magnetic field, the condensate field is concentrated near the surface of the “mathematical” cylinder, which implies a large influence. Thereby we may conclude that passive rotation of the frame and active rotation of the empty “physical” cylinder with a non-transparent wall represent, in a sense, different cases. Although in the former case interpretation causes questions, the latter case allows for a clearer physical interpretation. Thereby we focused on the latter case. Thus we related the metric to the rotating empty cylinder and considered the case when the flux tube (solenoid) of radius $R_{s}\leq R$ generating the uniform magnetic field for $r<R_{s}$, see eq. (52), is at rest in the frame associated with the rotating cylinder of radius $R$ (rotating frame). The aim was to describe the response of the scalar field vacuum to such a magnetic field in the rotating frame.
@@
In the foregoing section, to be specific, we considered two configurations of the magnetic field: an infinitely thin flux tube (the solenoid of radius $R_{s}\ll R$ generating uniform magnetic field inside it) and a field homogeneous in the whole cylinder (the solenoid of radius $R$). The case of the uniform magnetic field in the whole volume of the rotating cylinder was previously studied in [16, 17, 20]. The case of a thin flux tube in the rotating frame was not studied before, to the best of our knowledge. In both cases we found a numerical solution of the linearized equation, as an approximation, and a numerical solution of the full non-linear Gross-Pitaevskii-like (GP-like) equation, as well as found approximate analytical solutions, where possible. For the solution of the linearized equation, in the case of the infinitely thin flux tube, the condensate energy is expressed in terms of zeros of the Bessel function. For the case of the uniform magnetic field in the cylinder one needs to find a zero of the Kummer function, which problem is more delicate, especially for large values of parameters like the rotation frequency. However, for some special relations between magnetic flux and orbital momentum, e.g., when the magnetic flux $\delta_{\Phi}$ and the orbital momentum $l$ are connected as $\delta_{\Phi}=l+1$, an explicit solution exists. In the numerical approach, we adopted a relaxation method and a 4th order Runge-Kutta approach. Moreover, for the thin flux tube in the case $\delta_{\Phi}=l$ we found the approximate analytical solution of the GP-like equation. The latter case is of special interest, since we showed that the energy of the produced condensate takes a minimum precisely for $l=\delta_{\phi}$, see figure 3.